OpenStudy (anonymous):

use implicit differentiation to find the slope of the tangent line to the curve 3^x + log_2(xy) = 10 at the point (2,1) and use it to find the equation of the tangent line in the form y = mx + b

OpenStudy (anonymous):

im not really sure how to differentiate this if anyone could explain that to me

OpenStudy (experimentx):

your differeantiation is 3^3*ln(3) + 1/(xy)*{1+xdy/dx} = 0 find the value of dy/dx form here, put the value of m = dy/dx on slope, and put the value of x,y on the that eqn of tangent, find the value of b, you have your tangent

OpenStudy (anonymous):

i really don't think i know how to solve for dy/dx my professor just started implicit differentiation and I'm lost on this stuff

OpenStudy (experimentx):

okay differentiate this equation x^2+y=0

OpenStudy (anonymous):

2x+1=0

OpenStudy (experimentx):

no ... it's 2x+dy/dx = 0, here y is a variable that depends on x,

OpenStudy (experimentx):

of course you can see that, y = -x^2, which is a parabola, on this particular case, you can separate y, but this is not always the case

OpenStudy (anonymous):

ok so still how exactly do i get dy/dx by itself

OpenStudy (experimentx):

differentiate this wrt x y = -x^2

OpenStudy (experimentx):

you can differentiate y using chain rule, for eg you have x^2+y^2 = 25, try differentiating this

OpenStudy (anonymous):

2x+2y(dy/dx=0

OpenStudy (anonymous):

2x+2y(dy/dx)=0

OpenStudy (experimentx):

yes, so you have differential equation, though you don't know that is y, you can find slope of tangent at given point okay, what is slope of tangent at (0,5)

OpenStudy (anonymous):

the actual act of differentiating its figuring out how to separate the dy/dx to the other side in order to solve for it

OpenStudy (experimentx):

i wouldn't say solving it though, you can find the function for slope of tangent of given curve

OpenStudy (anonymous):

I'm sorry, I'm not sure what you mean

OpenStudy (experimentx):

f(x,y) = 0 gives a curve in x,y plane .. i guess you know that

OpenStudy (anonymous):

so I'm getting the slope as -10.05 when i plug in my x and y

OpenStudy (experimentx):

yeah, you get it because you know the FUNCTION OF SLOPE, and PUTTING x,y in FUNCTION GIVES the value of function, which is slope and also, you know that dy/dx gives the tangent ... right, so it must the the slope of tangent ... now you know it!!!